Assuming Raoult’s law applies, calculate the vapor pressure of a solution of 45 g of glucose (C6H12O6) in 95 g of water at 25 degrees Celsius


1
Expert's answer
2022-02-25T00:32:04-0500

Solution:

Сalculate the moles of each component of the solution:

The molar mass of glucose (C6H12O6) is 180 g/mol

The molar mass of water (H2O) is 18.015 g/mol

Therefore,

Moles of C6H12O6 = (45 g C6H12O6) × (1 mol C6H12O6 / 180 g C6H12O6) = 0.25 mol C6H12O6

Moles of H2O = (95 g H2O) × (1 mol H2O / 18.015 g H2O) = 5.27 mol H2O


Сalculate the mole fraction of each component of the solution:

Mole fraction (χ) = Moles / Total moles

Total moles of solution = Moles of C6H12O6 + Moles of H2O = 0.25 mol + 5.27 mol = 5.52 mol

Therefore,

χC6H12O6 = Moles of C6H12O6 / Total moles of solution = (0.25 mol) / (5.52 mol) = 0.0453

χH2O = Moles of H2O / Total moles of solution = (5.27 mol) / (5.52 mol) = 0.9547


Raoult’s law states that the vapor pressure of a solvent above a solution is equal to the vapor pressure of the pure solvent at the same temperature scaled by the mole fraction of the solvent present:

Psolution = χsolvent × Psolvent


The vapor pressure of water at room temperature (25°C) is 23.8 mmHg.

Therefore,

Psolution = 0.9547 × 23.8 mmHg = 22.72 mmHg

Psolution = 22.72 mmHg


Answer: The vapor pressure of a solution is 22.72 mmHg